### v1

# Influence of thermal noise on the field-driven dynamics of the non-collinear antiferromagnet Mn3Sn

Abstract
$\mathrm{Mn_{3}Sn}(0\overline{1}\overline{1}0)[0001]$ experiences a tensile strain when grown epitaxially on $\mathrm{MgO}(110)[001]$, and thus the energy landscape changes from six-fold symmetry to two-fold symmetry. External magnetic field further breaks the symmetry and the resulting energy landscape is sensitive to the field orientation relative to the easy axis. In the presence of thermal noise, the relaxation of the magnetic octupole moment in a strained Mn3Sn film is composed of four distinct escape processes involving the two saddle points and two equilibrium states in the energy landscape. Here, we apply harmonic transition-state theory to derive analytical expressions for the inter-well escape time and octupole moment relaxation time, both influenced by an external symmetry-breaking magnetic field and finite thermal noise in the intermediate-to-high damping regime. The analytical predictions are in strong agreement with comprehensive numerical simulations based on coupled LLG equations. The results presented here are crucial toward realizing Mn3Sn’s applications in random number generation and probabilistic computing.
preprint: AIP/123-QED
Non-collinear chiral antiferromagnets (AFMs) of the form Mn3X (X = Sn, Ge, Ir, and Pt) Shukla, Qian, and Rakheja (2025); Rimmler, Pal, and Parkin (2025) have gained significant attention owing to their large anomalous Hall effect (AHE), Nakatsuji, Kiyohara, and Higo (2015); Nayak et al. (2016); Chen, Niu, and MacDonald (2014) spin Hall effect (SHE), Zhang et al. (2016) anomalous Nernst effect (ANE), Ikhlas et al. (2017); Hong et al. (2020) magneto-optical Kerr effect (MOKE), Higo et al. (2018) and finite tunneling magnetoresistance (TMR). Dong et al. (2022); Chen et al. (2023); Qin et al. (2023); Wang et al. (2024) The anomalous Hall conductivity (AHC) of Mn3Sn, a prototypical non-collinear chiral AFM, can range from 30 to 40  $\Omega^{-1}\cdot$cm-1 at room temperature, which is attributed to the breaking of time-reversal symmetry (TRS). Liu et al. (2023); Yano et al. (2024) In its bulk form and below its Néel temperature ($T_{\mathrm{N}}$) of approximately 420 K, Mn3Sn exhibits a six-fold degenerate energy landscape. Higo et al. (2018); Sung et al. (2018) However, in an epitaxially grown Mn3Sn/MgO(110) [001] bilayer structure, the energy landscape is found to possess two-fold symmetry with the emergence of a perpendicular magnetic anisotropy (PMA) owing to an epitaxial tensile strain. Higo et al. (2022); Yoon et al. (2023)
Current-driven dynamics, including deterministic switching and chiral oscillations, have been experimentally explored in polycrystalline as well as epitaxial films of thicknesses ranging from sub-10 nm to 100 nm. Tsai et al. (2020); Takeuchi et al. (2021); Yan et al. (2022); Pal et al. (2022); Krishnaswamy et al. (2022); Higo et al. (2022); Yoon et al. (2023); Xu et al. (2023); Yoo et al. (2024); Zheng et al. (2025); Lee et al. (2025) The AHE signal in the case of polycrystalline films, measured in a typical Hall bar setup, is considered to be an average of the final magnetic states in the various composing grains. Ikeda et al. (2018); Xie et al. (2022) While in some cases pure spin-orbit torque (SOT) is considered to be the fundamental mechanism behind the observed dynamics, Tsai et al. (2020); Takeuchi et al. (2021); Higo et al. (2022); Yoon et al. (2023) others have attributed the observed changes in the AHE signal to a combination of Joule heating and SOT (i.e., demagnetization of the polycrystalline Mn3Sn film as the temperature rises above $T_{\mathrm{N}}$ due to Joule heating, followed by SOT-driven re-magnetization as the film cools below $T_{\mathrm{N}}$ during the slow decrease in current. Pal et al. (2022); Krishnaswamy et al. (2022); Yoo et al. (2024) The SOT-driven oscillation Zhao, He, and Cai (2021); Shukla and Rakheja (2022); Lund et al. (2023) and switching Shukla, Qian, and Rakheja (2023); Xu et al. (2024); Shukla, Qian, and Rakheja (2024); Zhao et al. (2025) dynamics in both six-fold and two-fold energy degenerate Mn3Sn crystals have also been explored theoretically. In particular, these works have extended the ferromagnetic theory to shed light on the threshold current requirement for initiating dynamics as well as the dependence of the various threshold currents on the material properties. They have also explored the possible dependence of the oscillation frequency and the switching time on the input stimuli and the material parameters. The manipulation of the order parameter in Mn3Sn via SOT, together with a finite TMR in all-Mn3Sn junctions, can enable the development of AFM-based spintronic devices that are fully compatible with electronic circuitry.
Despite significant progress in material synthesis and device functionality of $\mathrm{Mn_{3}Sn}$, studies investigating the impact of thermal noise on the dynamics and stability of the octupole moment in scaled Mn3Sn bits remain limited. Thermal noise is known to impose a lower bound on the bit error rate in ferromagnetic memories, Coffey and Kalmykov (2012); Kaiser et al. (2019) yet it can also be harnessed to generate random numbers Shukla et al. (2023); Chun et al. (2015) in hardware and potentially enable the paradigm of probabilistic computing. Shao et al. (2023) The thermal stability of Mn3Sn nanodots of diameters ranging from 175 to 1000 nm was experimentally investigated in Ref. [(42)]. The energy barrier for switching between the equilibrium states was extracted using the Néel-Arrhenius activation model, and it was found that the energy barrier decreased with the nanodot size below the nucleation diameter ($\sim 300$  nm for the sample). Kobayashi et al. Kobayashi et al. (2023) employed pulsed SOT measurements to extract the energy barrier by analyzing the relationship between the critical switching voltage and pulse duration. In Kobayashi’s sample, a robust AHE signal was measured in sub-100 nm grains, underscoring the promise of Mn3Sn for dense non-volatile memory applications.
Konakanchi et al. Konakanchi et al. (2025) presented a theoretical analysis of the relaxation time of the octupole order parameter, and its electrical tunability, in chiral AFMs. The relaxation of the octupole moment was found to depend either on the precessional dephasing when the energy barrier of the octupole was much smaller than the thermal energy, or given by the thermally activated escape over the barrier in the high-barrier regime.
Figure 1: (a) Kagome lattice of Mn3Sn. (b) Octupole moment and relevant angles in Cartesian coordinate system.
Konakanchi’s analysis indicates that the markedly shorter relaxation times of chiral AFMs—on the order of picoseconds—are driven by strong exchange interactions. However, their model neglects any external magnetic fields, which are typically incorporated in experimental studies and are essential for achieving more diverse manipulations of the order parameter.
In this paper, we investigate the impact of thermal noise on the stability of a strained $\mathrm{Mn_{3}Sn}$ bit in the presence of an external magnetic field. We present a comprehensive analysis of the mechanisms underlying the free-energy symmetry breaking, examining the dependence of free energy on both the magnitude and orientation of the applied field, as well as the associated fluctuation modes. The interwell escape time of the magnetic octupole moment is theoretically evaluated as a function of the external field and temperature, with results corroborated by numerical simulations in the low-energy barrier regime ($\Delta E_{0}\leq 5k_{\mathrm{B}}T$). We further extend our theoretical framework to encompass the relaxation dynamics of octupole moment and discuss prospective applications in random number generation and probabilistic computing.
Mn3Sn crystallizes into a hexagonal Kagome $D0_{19}$ lattice below its Néel temperature, exhibiting a slight excess of Mn atoms. Shukla, Qian, and Rakheja (2025) Sn atoms occupy the centers of the hexagons, whereas Mn atoms are located at the vertices. Within each Kagome layer, the Mn atoms contribute three magnetic moments that align at $120^{\circ}$ angles, forming a geometrically frustrated, noncollinear triangular spin structure. These moments are slightly canted toward the in-plane easy axis, giving rise to a non-zero net magnetization as shown in Fig. 1(a). The magnetic ground state typically exhibits six-fold degeneracy; Shukla, Qian, and Rakheja (2023) however, in the presence of an in-plane uniaxial strain, the symmetry is reduced to two-fold. Higo et al. (2022); Yoon et al. (2023)
The Hamiltonian of a single-domain $\mathrm{Mn_{3}Sn}$ is Tomiyoshi and Yamaguchi (1982); Park et al. (2018)
$\mathcal{H}(\mathbf{m})$ $=J_{\mathrm{E}}((1+\delta_{\mathrm{E}})\mathbf{m}_{1}\cdot\mathbf{m}_{2}+\mathbf{m}_{2}\cdot\mathbf{m}_{3}+\mathbf{m}_{3}\cdot\mathbf{m}_{1})+D_{\mathrm{M}}\mathbf{z}\cdot(\mathbf{m}_{1}\times\mathbf{m}_{2}+\mathbf{m}_{2}\times\mathbf{m}_{3}+\mathbf{m}_{3}\times\mathbf{m}_{1})-\sum_{i=1}^{3}{\left[K_{\mathrm{u}}(\mathbf{m}_{i}\cdot\mathbf{u}_{i})^{2}+M_{\mathrm{s}}\mathbf{H}_{\mathrm{a}}\cdot\mathbf{m}_{i})\right]},$ (1)
where $\mathbf{m}_{1}$, $\mathbf{m}_{2}$, and $\mathbf{m}_{3}$ are the magnetic moments of three sublattices, respectively. The parameters $M_{\mathrm{s}}$, $J_{\mathrm{E}}$, $D_{\mathrm{M}}$ and $K_{\mathrm{u}}$ represent the saturation magnetization of each sublattice, the symmetric exchange interaction constant, the asymmetric Dzyaloshinskii-Moriya Interaction (DMI) constant and the single-ion uniaxial magnetocrystalline anisotropy constant, respectively. In the Cartesian coordinates, $\mathbf{u}_{1}=-1/2\mathbf{x}+\sqrt{3}/2\mathbf{y}$, $\mathbf{u}_{2}=-1/2\mathbf{x}-\sqrt{3}/2\mathbf{y}$ and $\mathbf{u}_{3}=\mathbf{x}$ are the uniaxial easy axes of each sublattice. A positive (negative) $\delta_{E}$ corresponds to compressive (tensile) strain, enhancing (weakening) the exchange interaction among the corresponding sublattices. Unless otherwise noted, the values of material parameters used in our simulations are those listed in Table 1Yoon et al. (2023)
Using perturbation techniques, the Hamiltonian $\mathcal{H}_{\mathrm{oct}}$ of the single-domain Mn3Sn can be expressed in terms of $\varphi_{\mathrm{oct}}$ and $\theta_{\mathrm{oct}}$ (see Fig. 1(b)), which represent the azimuthal and polar angles of the octupole moment, respectively, as Yoon et al. (2023); He and Liu (2024); Shukla, Qian, and Rakheja (2024)
$\mathcal{H}(\theta_{\mathrm{oct}},\varphi_{\mathrm{oct}})=\frac{3}{4}M_{\mathrm{s}}H_{\mathrm{K}}\cos(2\varphi_{\mathrm{oct}})+\frac{3}{2}M_{\mathrm{s}}H_{\mathrm{J}}\cos^{2}\theta_{\mathrm{oct}}-\frac{M_{\mathrm{s}}H_{\mathrm{a}}\cos{(\theta_{\mathrm{H}})}}{J_{\mathrm{E}}+\sqrt{3}D_{\mathrm{M}}}\Big{[}K_{\mathrm{u}}\cos{(\varphi_{\mathrm{oct}}-\varphi_{\mathrm{H}})}+J_{\mathrm{E}}\delta_{E}\cos{(\varphi_{\mathrm{oct}}+\varphi_{\mathrm{H}})}\Big{]}.$ (2)
Here, we used $\mathbf{m}_{\mathrm{oct}}=\frac{1}{3}\mathcal{M}_{zx}\quantity[R\left(\frac{2\pi}{3}\right)\mathbf{m}_{1}+R\left(-\frac{2\pi}{3}\right)\mathbf{m}_{2}+\mathbf{m}_{3}]$ while $H_{\mathrm{J}}=\frac{(3J_{\mathrm{E}}+\sqrt{3}D_{\mathrm{M}})}{M_{s}}$ and $H_{\mathrm{K}}=-\frac{(4K_{\mathrm{u}}J_{\mathrm{E}}\delta_{\mathrm{E}})}{3M_{s}(J_{\mathrm{E}}+\sqrt{3}D_{\mathrm{M}})}$ are the strength of exchange field and uniaxial anisotropy field, respectively. $\theta_{\mathrm{H}}$ and $\varphi_{\mathrm{H}}$ are the polar and the azimuthal angle of the applied external field $\mathbf{H}_{\mathrm{a}}$, respectively.
Table 1: Material parameters of strained Mn3Sn thin film. Yoon et al. (2023)
Parameter Definition Value
$\alpha$ Damping coefficient 0.003
$J_{\mathrm{E}}~{}(\mathrm{J/m^{3}})$ Exchange constant $10^{8}$
$D_{\mathrm{M}}~{}(\mathrm{J/m^{3}})$ DMI constant $10^{7}$
$K_{\mathrm{u}}~{}(\mathrm{J/m^{3}})$ Uniaxial anisotropy constant $6.7\times 10^{5}$
$M_{\mathrm{s}}~{}(\mathrm{A/m})$ Saturation magnetization $4.46\times 10^{5}$
$\delta_{\mathrm{E}}$ Strain parameter $-2.6\times 10^{-4}$
$d_{\mathrm{z}}$ (nm) Thickness 8.8
Figure 2: Hamiltonian of $\mathrm{Mn_{3}Sn}$ as a function of $\varphi_{\mathrm{oct}}$ with $H_{\mathrm{a}}=0~{}\mathrm{T}$ and (a) $H_{\mathrm{ax}}=0.1~{}\mathrm{T}$ and (b) $H_{\mathrm{ay}}=0.1~{}\mathrm{T}$. Two equilibrium states of octupole moment are referred to as up state and down state in (a). (c) Splitting of the energy barriers as a function of the external magnetic field magnitude. $\Delta E_{0}$ denotes the energy barrier in absence of magnetic field.
From Eq. (2), strong exchange interactions constrain $\mathbf{m}_{\mathrm{oct}}$ to lie within the Kagome plane, resulting in two degenerate energy maxima at $\varphi_{\mathrm{oct}}=0$ and $\pi$, and two minima at $\varphi_{\mathrm{oct}}=\pi/2$ and $23\pi/$ 2, in the absence of an external field. The equilibrium states near $\varphi_{\mathrm{oct}}=\pi/2$ and $23\pi/$ 2 are hereafter referred to as the up and down states, respectively. An applied external field breaks the symmetry of the local extrema. As shown in Fig. 2(a), when $\mathbf{H}_{\mathrm{a}}$ is applied along the $-x$ direction (i.e., $\varphi_{\mathrm{H}}=\pi$), the degeneracy of the energy maxima is lifted. For a film of volume $\mathcal{V}$, the energy barriers, $\Delta E_{\mathrm{x},1}$ and $\Delta E_{\mathrm{x},2}$, evaluated at $\varphi_{\mathrm{oct}}=0$ and $\pi$, respectively, are given as
$\Delta E_{\mathrm{x,1(2)}}=\mathcal{V}\Big{[}\frac{3}{2}M_{\mathrm{s}}H_{\mathrm{K}}\pm 3M_{\mathrm{s}}H_{\mathrm{x,eff}}+\frac{3}{2}M_{\mathrm{s}}\frac{H_{\mathrm{x,eff}}^{2}}{H_{\mathrm{K}}}\Big{]},$ (3)
where $H_{\mathrm{x,eff}}=H_{\mathrm{ax}}\frac{(K_{\mathrm{u}}+J_{\mathrm{E}}\delta_{\mathrm{E}})}{3(J_{\mathrm{E}}+\sqrt{3}D_{\mathrm{M}})}$ is the normalized external field applied along the $-x$ direction. In this configuration, the up and down states remain energetically equivalent; however, thermal fluctuations preferentially promote transitions across $\Delta E_{\mathrm{x,2}}$ due to its lower energy barrier.
In contrast, when the external field is applied along the $-y$ direction, it primarily perturbs the energy minima, as shown in Fig. 2(b). The energy maxima remain degenerate, resulting in symmetric energy barriers on either side of each state. The analytical expressions for the energy barriers $\Delta E_{\mathrm{y,1}}$ and $\Delta E_{\mathrm{y,2}}$, corresponding to the up and down states, respectively, are
$\Delta E_{\mathrm{y,1(2)}}=\mathcal{V}\Big{[}\frac{3}{2}M_{\mathrm{s}}H_{\mathrm{K}}\mp 3M_{\mathrm{s}}H_{\mathrm{y,eff}}+\frac{3}{2}M_{\mathrm{s}}\frac{H_{\mathrm{y,eff}}^{2}}{H_{\mathrm{K}}}\Big{]},$ (4)
where $H_{\mathrm{y,eff}}=H_{\mathrm{ay}}\frac{(K_{\mathrm{u}}-J_{\mathrm{E}}\delta_{\mathrm{E}})}{3(J_{\mathrm{E}}+\sqrt{3}D_{\mathrm{M}})}$ denotes the normalized external field applied along the $-y$ direction. For comparison, Fig. 2(c) quantitatively illustrates the quadratic dependence of the energy barrier density on the magnitude of the external field. Additionally, we define $\Delta E_{0}$ as the energy barrier height in the absence of an external field, which serves as a reference in subsequent analyses.
Based on the harmonic transition-state theory (HTST), Coffey and Kalmykov (2012); Morse and Feshbach (1946) the escape time with intermediate-to-high (IHD) damping is formulated per Eq. (5). In this equation, $\varepsilon_{j,\mathrm{min(sp)}}$ denotes the $j$-th eigenvalue of the Hessian matrix in the harmonic approximation near the energy minimum (saddle point), and $\lambda_{+}$ is the positive eigenvalue of the linearized Laudau-Lifshitz-Gilbert (LLG) equation at the saddle point. $P_{\mathrm{min(sp)}}$ represents the number of Goldstone modes at the minimum (saddle point), and $V_{\mathrm{min(sp)}}$ denotes the corresponding phase-space volume. Additionally, $A(\delta E/k_{\mathrm{B}}T)$ is the depopulation factor that accounts for the coupling strength to the thermal bath, where $\delta E$ is the energy dissipated along an iso-energy contour encompassing the energy maximum. See Eq. (6) for $A(x)$.
In the theoretical analysis of escape time, we employ the angular variables $\theta_{\mathrm{oct}}$ and $\varphi_{\mathrm{oct}}$ to describe the dynamics of the octupole moment $\mathbf{m}_{\mathrm{oct}}$ within a perturbative framework. He and Liu (2024) Further methodological details are provided in the Supplementary Material . The escape time due to thermal fluctuations over a single saddle point, from the up state to the down state under an external field $H_{\mathrm{ax}}$ or $H_{\mathrm{ay}}$ is given in Eq. (7), where $k=1$ and $k=2$ correspond to the magnetic octupole crossing the saddle point near $\varphi_{\mathrm{oct}}=0$ and $\pi$, respectively. The parameters $\beta_{x}$ and $\beta_{y}$ in Eq. (7) are given as $\left((-1)^{k}\frac{H_{\mathrm{x,eff}}}{H_{\mathrm{K}}}-1\right)$ and $\left(3\frac{H_{\mathrm{y,eff}}^{2}}{H_{\mathrm{K}}^{2}}-1\right)$, respectively. Our calculations suggest that for the magnetic fields and energy barriers considered in this work, $\quantity[A(\delta E/k_{\mathrm{B}}T)]^{-1}$ can be approximated to unity with less than 5% error (see Supplementary Material for details).
$\tau_{\mathrm{esc}}=\left[A\left(\frac{\delta E}{k_{\mathrm{B}}T}\right)\right]^{-1}\frac{2\pi}{\lambda_{+}}\frac{V_{\mathrm{min}}}{V_{\mathrm{sp}}}(2\pi k_{\mathrm{B}}T)^{\frac{P_{\mathrm{sp}}-P_{\mathrm{min}}}{2}}\sqrt{\frac{\Pi^{\prime}_{j}|\varepsilon_{j,\mathrm{,sp}}|}{\Pi^{\prime}_{j}\varepsilon_{j,\mathrm{min}}}}e^{\frac{\Delta E}{k_{\mathrm{B}}T}}.$ (5)
[…]
$\tau_{\mathrm{escx,k}}=\quantity[A\quantity(\frac{\delta E}{k_{\mathrm{B}}T})]^{-1}\frac{4\pi(1+\alpha^{2})\exp(\Delta E_{\mathrm{x,k}}/k_{\mathrm{B}}T)}{\gamma\sqrt{|H_{\mathrm{x,eff}}-(-1)^{k+1}H_{\mathrm{K}}|}}\times\frac{\sqrt{H_{\mathrm{K}}}}{-\alpha\left[H_{J}+\beta_{x}H_{K}\right]+\sqrt{\alpha^{2}\left(H_{\mathrm{J}}-\beta_{x}H_{\mathrm{K}}\right)^{2}-4\beta_{x}H_{\mathrm{J}}H_{\mathrm{K}}}}.$ (7a)
$\tau_{\mathrm{escy,k}}=\quantity[A\quantity(\frac{\delta E}{k_{\mathrm{B}}T})]^{-1}\frac{4\pi(1+\alpha^{2})\exp(\Delta E_{\mathrm{y,1}}/k_{\mathrm{B}}T)}{\gamma H_{K}}\times\frac{\sqrt{|H_{\mathrm{y,eff}}^{2}-H_{K}^{2}|}}{-\alpha\left[H_{J}+\beta_{y}H_{K}\right]+\sqrt{\alpha^{2}\left(H_{J}-\beta_{y}H_{K}\right)^{2}-4\beta_{y}H_{J}H_{K}}}.$ (7b)
$\frac{1}{\tau_{\mathrm{escx(y),}\uparrow}}=\frac{1}{\tau_{\mathrm{escx(y),1}}}+\frac{1}{\tau_{\mathrm{escx(y),2}}}.$ (8)
$m_{\mathrm{oct,y}}(t)$ $=\frac{\tau_{\mathrm{esc,}\uparrow}-\tau_{\mathrm{esc,}\downarrow}}{\tau_{\mathrm{esc,}\uparrow}+\tau_{\mathrm{esc,}\downarrow}}+2\left[m_{\mathrm{oct,y}}(0)-\frac{\tau_{\mathrm{esc,}\uparrow}}{\tau_{\mathrm{esc,}\uparrow}+\tau_{\mathrm{esc,}\downarrow}}\right]\exp\quantity(-\frac{\tau_{\mathrm{esc,}\uparrow}+\tau_{\mathrm{esc,}\downarrow}}{\tau_{\mathrm{esc,}\uparrow}\tau_{\mathrm{esc,}\downarrow}})t.$ (10)
As indicated by Eq. (7), under an external field applied along the $y$-axis, the escape times of the magnetic octupole in the up state over the two saddle points to the down state are identical due to the preserved symmetry of the energy maxima. In contrast, an external field applied along the $x$-axis introduces asymmetry between the energy maxima, resulting in different escape times for the magnetic octupole over the two barriers. For the magnetic octupole in the down state, the escape time over the two barriers to the up state can be evaluated using Eq. 7(a), when the magnetic field is applied along the $x$-axis. However, for the magnetic field along the $y$-axis, $\Delta E_{\mathrm{y,1}}$ in Eq. 7(b) should be replaced by $\Delta E_{\mathrm{y,2}}$ for an accurate definition of the escape time. Assuming the transitions from the up (down) state to the down (up) state over the two barriers to be independent, the total escape time satisfies the relationship in Eq. (8).
Figure 3: The schematic of escape time from up state to down state while numerical extracted data is in markers and analytical predictions are plotted with lines. $\Delta E_{0}$ ranges from $k_{\mathrm{B}}T$ to $5k_{\mathrm{B}}T$ and field is applied along (a) $x$-axis and (b) $y$-axis. (c) The time evolution of octupole moment with external field and $\Delta E_{0}=4k_{\mathrm{B}}T$.
To validate Eqs. (7) and (8), the dynamics of the magnetic octupole is numerically explored via three coupled stochastic LLG equations, Mayergoyz, Bertotti, and Serpico (2009); Go et al. (2022) for $\Delta E_{0}$ ranging from $k_{\mathrm{B}}T$ to $5k_{\mathrm{B}}T$ and $H_{\mathrm{a}}\leq 0.1~{}\mathrm{T}$. For each sublattice, $i$ (= 1,2,3), the LLG equation is given as $\partialderivative{\mathbf{m}_{i}}{t}=-\gamma(\mathbf{m}_{i}\crossproduct\mathbf{H}_{i}^{\mathrm{eff}})+\alpha\quantity(\mathbf{m}_{i}\times\partialderivative{\mathbf{m}_{i}}{t})$, where $\mathbf{H}_{i}^{\mathrm{eff}}=-\frac{1}{M_{\mathrm{s}}}\frac{\partial\mathcal{H}}{\partial\mathbf{m}_{i}}+\mathbf{H}_{\mathrm{thermal},i}$ is the effective magnetic field and includes the thermal noise field, $\mathbf{H}_{\mathrm{thermal},i}=\sqrt{\frac{2\alpha k_{\mathrm{B}}T}{\gamma\mu_{0}M_{\mathrm{s}}^{2}\mathcal{V}\Delta t}}\bm{\xi}^{i}_{t}$. In this work, we assume $\mathbf{H}_{\mathrm{thermal},i}$ to be Gaussian with zero mean as well as uncorrelated in space, time, and between sublattices, such that $\bm{\xi}^{i}_{t}\sim\mathcal{N}(0,1)$Go et al. (2022); Ament et al. (2016) Here, $\gamma$ and $\mu_{0}$ denote the gyromagnetic ratio and vacuum permeability, respectively, while $\Delta t$ is the time step used in the simulation. Other parameters are defined in Table 1.
In Figs. 3(a) and 3(b), the numerically computed escape time from up state to down state as a function of $H_{\mathrm{ax}}$ and $H_{\mathrm{ay}}$, for five different $\Delta E_{0}$, shown using symbols, is compared to the analytic predictions of Eq. (7) and (8), represented by solid lines. We find that, for both directions of the applied field, the analytical predictions agree very well with the numerical data for $\Delta E_{0}\geq 2k_{\mathrm{B}}T$, though a small discrepancy emerges at higher fields ($H_{\mathrm{a}}\gtrsim 75~{}\mathrm{mT}$) and increases with both $H_{\mathrm{ax}}$ and $H_{\mathrm{ay}}$. This is due to a decrease in the effective barrier heights ($\Delta E_{\mathrm{x,1}}$ and $\Delta E_{y,1}$) below the valid limits of the HTST theory. In the case of $\Delta E_{0}=k_{\mathrm{B}}T$ the discrepancy between the analytic model and numerical data begins at a comparatively lower field ($H_{\mathrm{a}}\gtrsim 45~{}\mathrm{mT}$). For the two directions of magnetic fields, the escape time appears to be more sensitive to $H_{\mathrm{ay}}$ in the low-field regime. This is because an increase in $H_{\mathrm{ay}}$ decreases $\Delta E_{\mathrm{y,1}}$, favoring the octupole transitions from the up state to the down state over both the saddle points equally, thereby speeding up the process. On the other hand, when $H_{\mathrm{ax}}$ increases above zero, thermal field-driven transitions from the up state to the down state are comparatively slower as only $\Delta E_{\mathrm{x,2}}$ decreases but $\Delta E_{\mathrm{x,1}}$ increases. Although not shown here, we find $\tau_{\mathrm{escx,\downarrow}}$ to be same as the results shown in Fig. 3(a), but $\tau_{\mathrm{escy,\downarrow}}$ increases exponentially with $H_{\mathrm{ay}}$ due to increase in $\Delta E_{\mathrm{y,2}}$ (see Supplementary Material).
The thermal relaxation in Mn3Sn arises from the collective dynamics of multiple magnetic moments transitioning back and forth between two energy wells. The temporal evolution of the spin population is governed by the rate equations Brown Jr (1963); Brown (1979) specified in Eq. (9), where $n_{\uparrow}$ and $n_{\downarrow}$ denote the number of spins in the up and down states, respectively. Given the constraint $n_{\uparrow}+n_{\downarrow}=N_{\mathrm{total}}$ and the definition of the octupole moment’s $y$ component as $m_{\mathrm{oct,y}}=\frac{(n_{\uparrow}-n_{\downarrow})}{N_{\mathrm{total}}}$, the time evolution of $m_{\mathrm{oct,y}}$ is given in Eq. (10), where $m_{\mathrm{oct,y}}(0)$ denotes the initial octupole moment. From this expression, the thermal equilibrium state $m_{\mathrm{oct,y}}(\infty)$ is given by $m_{\mathrm{oct,y}}(\infty)=\frac{\tau_{\mathrm{esc,}\uparrow}-\tau_{\mathrm{esc,}\downarrow}}{\tau_{\mathrm{esc,}\uparrow}+\tau_{\mathrm{esc,}\downarrow}}.$ In particular, when the external field is applied along the $x$-axis, the escape times become equal, $\tau_{\mathrm{esc,}\uparrow}=\tau_{\mathrm{esc,}\downarrow}$, leading to a vanishing steady-state magnetization, i.e., $m_{\mathrm{oct,y}}(\infty)=0$. On the other hand, as the external field along the $y$-axis increases, thermal transitions in one direction become more favorable than the reverse process, leading to a well defined steady-state magnetization, i.e., $m_{\mathrm{oct,y}}(\infty)\to\pm 1$. The relaxation time associated with this two-state system is $\tau_{\mathrm{relax}}=\frac{\tau_{\mathrm{esc,}\uparrow}\,\tau_{\mathrm{esc,}\downarrow}}{\tau_{\mathrm{esc,}\uparrow}+\tau_{\mathrm{esc,}\downarrow}}$, which characterizes the timescale over which $m_{\mathrm{oct,y}}$ decays to $1/e$ of its initial deviation from equilibrium. The temporal evolution of $m_{\mathrm{oct,y}}$ for $\Delta E_{0}=4k_{\mathrm{B}}T$, as shown in Fig. 3(c), illustrates the dependence of the relaxation time and steady-state magnetization on the magnitude and direction of the external field.
Figure 4: (a) The ACF as a function of sample interval with $x$-axis external field. (b) The probability of finding the octupole moment in the down state at steady state with a sweep of $H_{\mathrm{ay}}$ and pulse width.
The distinct thermal fluctuation mechanisms provide valuable insights into the prospective applications of $\mathrm{Mn_{3}Sn}$. When an external field is applied along the $x$-axis, the two equilibrium states remain symmetric, resulting in equal occupation probabilities. This inherent symmetry renders the system a promising candidate for random number generation. As shown in Fig. 4(a), for a single $\mathrm{Mn_{3}Sn}$-based magnetic tunnel junction, the autocorrelation function (ACF) of the octupole moment decays to zero with a 2-ns sampling interval. Moreover, increasing the magnitude of the applied field effectively reduces the required sampling interval, whereas the typical sampling interval for ferromagnetic random number generators is on the order of $\mu s$ or even higher. Wang et al. (2022); Chen, Zhang, and Xiao (2022) In contrast, when the external field is applied along the $y$-axis, the symmetry between the two equilibrium states is broken, enabling probabilistic control of the octupole moment $\mathbf{m}_{\mathrm{oct}}$. Figure 4(b) presents the temporal evolution of the probability of occupying the down state under various field magnitudes, thereby confirming the effectiveness of $H_{\mathrm{ay}}$ in modulating the switching probability.
In conclusion, we comprehensively demonstrated the external field-assisted thermal behaviors of $\mathrm{Mn_{3}Sn}$ within the energy barrier $\leq 5k_{\mathrm{B}}T$. The energy landscape symmetry breakdown is dependent on the field orientation, in which case the theoretical prediction of escape time varies and contributes to two different relaxation modes. The two modes provide prospective applications in random number generation and probabilistic computing while the frequency can reach above GHz. This work focuses on low barrier magnets ($\leq 5k_{\mathrm{B}}T$) and low external field ($\leq 0.1~{}\mathrm{T}$), which is the effective regime of perturbation theory. However, current experiments are reported within high barrier regime ($\geq 40k_{\mathrm{B}}T$) and the switching field is up to $0.5~{}\mathrm{T}$ so our future work will explore the thermal behaviors of $\mathrm{Mn_{3}Sn}$ within this regime. Our work highlights the relevance of thermal noise in establishing the physical understanding and operating principles of scaled $\mathrm{Mn_{3}Sn}$ devices and enable applications such as random number generation and probabilistic computing.
Supplementary Material
See the supplementary material for a derivation of the escape time of the octupole, numerical LLG simulation, and regimes of model validity.
Acknowledgment
This research was supported by the NSF through the University of Illinois at Urbana-Champaign Materials Research Science, Engineering Center DMR-1720633.
Author Declarations
Conflict of Interest
The authors have no conflicts to disclose.
Author Contributions
Siyuan Qian: Conceptualization (equal); Data curation (lead); Formal analysis (lead); Investigation (lead); Methodology (lead); Software (lead); Validation (lead); Writing – original draft (lead); Writing – review & editing (supporting). Ankit Shukla: Data curation (supporting); Formal analysis (supporting); Software (supporting), Writing – review & editing (lead). Shaloo Rakheja: Conceptualization (lead); Funding acquisition (lead); Project administration (lead); Resources (lead); Supervision (lead); Writing – original draft (supporting); Writing – review & editing (supporting).
Data Availability
The data that support the findings of this study are available from the corresponding author upon reasonable request.
Supplementary Material
.1 The Landau-Lifshitz-Gilbert (LLG) equation
The LLG equation for a magnetic moment vector $(\hat{r},\hat{\theta},\hat{\varphi})$ in the spherical coordinates is given as
$\dot{\theta}$ $=\frac{\gamma}{1+\alpha^{2}}\left(H_{\varphi}+\alpha H_{\theta}\right),$ (11a)
$\dot{\varphi}$ $=\frac{\gamma}{1+\alpha^{2}}\left(-H_{\theta}+\alpha H_{\varphi}\right)\csc\theta,$ (11b)
where $|\hat{r}|\equiv 1$ and the effective magnetic field in spherical coordinate is obtained from the Hamiltonian, $\mathcal{H}$, as
$H_{\varphi}$ $=-\frac{1}{M_{\mathrm{s}}}\frac{\partial\mathcal{H}}{\partial\varphi},$ (12a)
$H_{\theta}$ $=-\frac{1}{M_{\mathrm{s}}\sin\theta}\frac{\partial\mathcal{H}}{\partial\theta}.$ (12b)
In these equations, $\gamma$, $\alpha$, and $M_{\mathrm{s}}$ represent the gyromagnetic ratio, damping coefficient, and saturation magnetization, respectively, of the magnetic material.
In the case of $\mathrm{Mn_{3}Sn}$, we model the dynamics of the magnetic octupole moment, which is defined as
$\mathbf{m}_{\mathrm{oct}}=\frac{1}{3}\mathcal{M}_{zx}\left[R\left(\frac{2\pi}{3}\right)\mathbf{m}_{1}+R\left(-\frac{2\pi}{3}\right)\mathbf{m}_{2}+\mathbf{m}_{3}\right],$ (13)
where $\mathcal{M}_{zx}$ is the mirror operation against $zx$ plane and $\mathcal{R}$ is the rotation operation in $xy$ plane, so $\varphi$ in Eq. (11) is replace by $-\varphi$. Owing to the strong exchange interaction between the sublattice vectors, $\mathbf{m}_{i}$ (with $i=1,2,3$), $\mathrm{\theta}_{\mathrm{oct}}$ is restricted around $\pi/2$ and therefore, LLG equation and the effective field for $\mathbf{m}_{\mathrm{oct}}$ is
$\dot{\theta}_{\mathrm{oct}}$ $=\frac{\gamma}{1+\alpha^{2}}\left(H_{\varphi}(-\varphi_{\mathrm{oct}})+\alpha H_{\theta}\right),$ (14a)
$\dot{\varphi}_{\mathrm{oct}}$ $=\frac{\gamma}{1+\alpha^{2}}\left(-H_{\theta}+\alpha H_{\varphi}(-\varphi_{\mathrm{oct}})\right).$ (14b)
To average the joint contribution of the three sublattices, the effective field is scaled by a factor of $\frac{1}{3}$e and Liu (2024); Konakanchi et al. (2025) leading to
$H_{\varphi}$ $=-\frac{1}{3M_{\mathrm{s}}}\frac{\partial\mathcal{H}}{\partial\varphi_{\mathrm{oct}}},$ (15a)
$H_{\theta}$ $=-\frac{1}{3M_{\mathrm{s}}\sin\theta_{\mathrm{oct}}}\frac{\partial\mathcal{H}}{\partial\theta_{\mathrm{oct}}}.$ (15b)
Further simplification yields
$H_{\varphi}$ $=-\frac{1}{3M_{\mathrm{s}}}\frac{\partial\mathcal{H}_{\mathrm{oct}}}{\partial\varphi_{\mathrm{oct}}}=\frac{1}{2}H_{\mathrm{K}}\sin(2\varphi_{\mathrm{oct}})-\frac{H_{\mathrm{a}}\cos{(\theta_{\mathrm{H}})}}{3(J_{\mathrm{E}}+\sqrt{3}D_{\mathrm{M}})}\Big{[}K_{\mathrm{u}}\sin{(\varphi_{\mathrm{oct}}-\varphi_{\mathrm{H}})}+J_{\mathrm{E}}\delta_{E}\sin{(\varphi_{\mathrm{oct}}+\varphi_{\mathrm{H}})}\Big{]},$ (16a)
$H_{\theta}$ $=-\frac{1}{3M_{\mathrm{s}}}\frac{\partial\mathcal{H}_{\mathrm{oct}}}{\partial\theta_{\mathrm{oct}}}=H_{\mathrm{J}}\cos{\theta_{\mathrm{oct}}}\sin{\theta_{\mathrm{oct}}}.$ ) . (16b)
.2 Field-assisted switching
.2.1 $H_{\mathrm{ay}}$-assisted dynamics
  • Eigenvalue of linearized LLG
With $\varphi_{\mathrm{H}}=3\pi/2$, the Hamiltonian is
$\mathcal{H}_{\mathrm{oct}}=\frac{3}{4}M_{\mathrm{s}}H_{\mathrm{K}}\cos(2\varphi_{\mathrm{oct}})+\frac{3}{2}M_{\mathrm{s}}H_{\mathrm{J}}\cos^{2}\theta_{\mathrm{oct}}+3M_{\mathrm{s}}H_{\mathrm{y,eff}}\sin\varphi_{\mathrm{oct}},$ (17)
in which case local minimum is at $\varphi_{\mathrm{oct}}=\pi/2$ and $23\pi/$ 2, while local maximum is at $\sin\varphi_{\mathrm{oct}}=\frac{H_{\mathrm{y,eff}}}{H_{\mathrm{K}}}$. Using Eq. (17) in Eq. (14), we get
$\dot{\theta}_{\mathrm{oct}}$ $=\frac{\gamma}{1+\alpha^{2}}\left(H_{\mathrm{K}}\sin\varphi_{\mathrm{oct}}\cos\varphi_{\mathrm{oct}}-H_{\mathrm{y,eff}}\cos\varphi_{\mathrm{oct}}+\alpha H_{J}\cos\theta_{\mathrm{oct}}\sin\theta_{\mathrm{oct}}\right),$ (18a)
$\dot{\varphi}_{\mathrm{oct}}$ $=\frac{\gamma}{1+\alpha^{2}}\left(-H_{J}\cos\theta_{\mathrm{oct}}\sin\theta_{\mathrm{oct}}+\alpha H_{\mathrm{K}}\sin\varphi_{\mathrm{oct}}\cos\varphi_{\mathrm{oct}}-\alpha H_{\mathrm{y,eff}}\cos\varphi_{\mathrm{oct}}\right).$ (18b)
The linearized LLG equation of $\mathbf{m}_{\mathrm{oct}}$ at the saddle point $\theta_{\mathrm{oct}}=\pi/2$ and $\sin\varphi_{\mathrm{oct}}=H_{\mathrm{y,eff}}/H_{\mathrm{K}}$ is
$\begin{pmatrix}\delta\dot{\theta}_{\mathrm{oct}}\\ \delta\dot{\varphi}_{\mathrm{oct}}\end{pmatrix}=\frac{\gamma}{1+\alpha^{2}}\begin{pmatrix}-\alpha H_{J}&\beta_{y}H_{K}\\ -H_{J}&-\alpha\beta_{y}H_{K}\end{pmatrix}\begin{pmatrix}\delta\dot{\theta}_{\mathrm{oct}}\\ \delta\dot{\varphi}_{\mathrm{oct}}\end{pmatrix},$ (19)
where $\beta_{y}=3\frac{H_{\mathrm{y,eff}}^{2}}{H_{\mathrm{K}}^{2}}-1$ and the positive eigenvalue of the matrix is
$\lambda_{+}=\frac{\gamma}{(1+\alpha^{2})}\frac{-\alpha\left[H_{\mathrm{J}}+\beta_{y}H_{\mathrm{K}}\right]+\sqrt{\alpha^{2}\left(H_{\mathrm{J}}-\beta_{y}H_{\mathrm{K}}\right)^{2}-4\beta_{y}H_{\mathrm{J}}H_{\mathrm{K}}}}{2}.$ (20)
  • Eigenvalue of Hessian matrix
The eigenvalues of the Hessian matrix at the minimum and saddle points(sp) are
$\varepsilon_{\mathrm{1,min}}$ $=\mathcal{V}\frac{\partial^{2}\mathcal{H}_{\mathrm{oct}}}{\partial\theta_{\mathrm{oct}}^{2}}\Big{|}_{\theta_{\mathrm{oct}}=\pi/2,\varphi_{\mathrm{oct}}=\pi/2}=3M_{s}H_{J}\mathcal{V},$ (21a)
$\varepsilon_{\mathrm{2,min}}$ $=\mathcal{V}\frac{\partial^{2}\mathcal{H}_{\mathrm{oct}}}{\partial\theta_{\mathrm{oct}}^{2}}\Big{|}_{\theta_{\mathrm{oct}}=\pi/2,\varphi_{\mathrm{oct}}=3\pi/2}=3M_{s}H_{J}\mathcal{V},$ (21b)
$\varepsilon_{\mathrm{1,sp}}$ $=\mathcal{V}\frac{\partial^{2}\mathcal{H}_{\mathrm{oct}}}{\partial\theta_{\mathrm{oct}}^{2}}\Big{|}_{\theta_{\mathrm{oct}}=\pi/2,\varphi_{\mathrm{oct}}=\sin^{-1}\quantity(\frac{H_{\mathrm{y,eff}}}{H_{K}})}=3M_{s}H_{J}\mathcal{V},$ (21c)
$\varepsilon_{\mathrm{2,sp}}$ $=\mathcal{V}\frac{\partial^{2}\mathcal{H}_{\mathrm{oct}}}{\partial\theta_{\mathrm{oct}}^{2}}\Big{|}_{\theta_{\mathrm{oct}}=\pi/2,\varphi_{\mathrm{oct}}=\pi-\sin^{-1}\quantity(\frac{H_{\mathrm{y,eff}}}{H_{K}})}=3M_{s}H_{J}\mathcal{V},$ (21d)
$\varepsilon_{\mathrm{3,min}}$ $=\mathcal{V}\frac{\partial^{2}\mathcal{H}_{\mathrm{oct}}}{\partial\varphi_{\mathrm{oct}}^{2}}\Big{|}_{\theta_{\mathrm{oct}}=\pi/2,\varphi_{\mathrm{oct}}=\pi/2}=3M_{s}\mathcal{V}(H_{K}-H_{\mathrm{y,eff}}),$ (21e)
$\varepsilon_{\mathrm{4,min}}$ $=\mathcal{V}\frac{\partial^{2}\mathcal{H}_{\mathrm{oct}}}{\partial\varphi_{\mathrm{oct}}^{2}}\Big{|}_{\theta_{\mathrm{oct}}=\pi/2,\varphi_{\mathrm{oct}}=3\pi/2}=3M_{s}\mathcal{V}(H_{K}+H_{\mathrm{y,eff}}),$ (21f)
$\varepsilon_{\mathrm{3,sp}}$ $=\mathcal{V}\frac{\partial^{2}\mathcal{H}_{\mathrm{oct}}}{\partial\varphi_{\mathrm{oct}}^{2}}\Big{|}_{\theta_{\mathrm{oct}}=\pi/2,\varphi_{\mathrm{oct}}=\sin^{-1}\quantity(\frac{H_{\mathrm{y,eff}}}{H_{K}})}=3M_{s}\mathcal{V}\quantity(\frac{H_{\mathrm{y,eff}}^{2}}{H_{K}}-H_{K}),$ (21g)
$\varepsilon_{\mathrm{4,sp}}$ $=\mathcal{V}\frac{\partial^{2}\mathcal{H}_{\mathrm{oct}}}{\partial\varphi_{\mathrm{oct}}^{2}}\Big{|}_{\theta_{\mathrm{oct}}=\pi/2,\varphi_{\mathrm{oct}}=\pi-\sin^{-1}\quantity(\frac{H_{\mathrm{y,eff}}}{H_{K}})}=3M_{s}\mathcal{V}\quantity(\frac{H_{\mathrm{y,eff}}^{2}}{H_{K}}-H_{K}),$ (21h)
where $\mathcal{V}$ is the sample volume, and, therefore, we have
$\sqrt{\frac{\prod_{j}\left|\varepsilon_{j,\text{sp}}\right|}{\prod_{j}\varepsilon_{j,\min}}}=\sqrt{\frac{\quantity|\quantity(\frac{H_{\mathrm{y,eff}}}{H_{K}})^{2}-1|^{2}}{\quantity|1-\quantity(\frac{H_{\mathrm{y,eff}}}{H_{K}})^{2}|}}=\sqrt{\quantity|\frac{H_{\mathrm{y,eff}}^{2}-H_{K}^{2}}{H_{K}^{2}}|}.$ (22)
In this work we only analyze $H_{\mathrm{y,eff}}<H_{\mathrm{K}}$ so there is no zero eigenvalue, which indicates no Goldstone mode Rózsa et al. (2019) in this system. So we have $P_{\mathrm{sp}}=P_{\mathrm{min}}=0$ and $V_{\mathrm{min}}/V_{\mathrm{sp}}=1$.
Finally, the escape time with external field applied along $-y$ axis can be evaluated as
$\tau_{\mathrm{escy}}$ $=\quantity[A\quantity(\frac{\delta E}{k_{\mathrm{B}}T})]^{-1}\frac{2\pi}{\lambda_{+}}\sqrt{\frac{\Pi^{\prime}_{j}|\varepsilon_{j,\mathrm{,sp}}|}{\Pi^{\prime}_{j}\varepsilon_{j,\mathrm{min}}}}e^{\frac{\Delta E}{k_{\mathrm{B}}T}}$ (23)
$=\quantity[A\quantity(\frac{\delta E}{k_{\mathrm{B}}T})]^{-1}\frac{4\pi(1+\alpha^{2})\exp(\frac{\Delta E_{\mathrm{y,1}}}{k_{\mathrm{B}}T})\sqrt{\quantity|H_{\mathrm{y,eff}}^{2}-H_{K}^{2}|}}{\gamma H_{K}\quantity(-\alpha\left[H_{J}+\beta_{y}H_{K}\right]+\sqrt{\alpha^{2}\left(H_{J}-\beta_{y}H_{K}\right)^{2}-4\beta_{y}H_{J}H_{K}})}.$
.2.2 $H_{\mathrm{ax}}$-assisted dynamics
  • Eigenvalue of linearized LLG
With $\varphi_{\mathrm{H}}=\pi$, the Hamiltonian is
$\mathcal{H}_{\mathrm{oct}}=\frac{3}{4}M_{\mathrm{s}}H_{\mathrm{K}}\cos(2\varphi_{\mathrm{oct}})+\frac{3}{2}M_{\mathrm{s}}H_{\mathrm{J}}\cos^{2}\theta_{\mathrm{oct}}+3M_{\mathrm{s}}H_{\mathrm{x,eff}}\cos\varphi_{\mathrm{oct}}.$ (24)
In this case, the LLG in spherical coordinates is given as
$\dot{\theta}_{\mathrm{oct}}$ $=\frac{\gamma}{1+\alpha^{2}}(H_{\mathrm{K}}\sin\varphi_{\mathrm{oct}}\cos\varphi_{\mathrm{oct}}+H_{\mathrm{x,eff}}\sin\varphi_{\mathrm{oct}}+\alpha H_{\mathrm{J}}\cos\theta_{\mathrm{oct}}\sin\theta_{\mathrm{oct}}),$ (25a)
$\dot{\varphi}_{\mathrm{oct}}$ $=\frac{\gamma}{1+\alpha^{2}}(-H_{\mathrm{J}}\cos\theta_{\mathrm{oct}}\sin\theta_{\mathrm{oct}}+\alpha H_{\mathrm{K}}\sin\varphi_{\mathrm{oct}}\cos\varphi_{\mathrm{oct}}+\alpha H_{\mathrm{x,eff}}\sin\varphi_{\mathrm{oct}}).$ (25b)
The saddle point corresponding to $\varphi_{\mathrm{oct}}=0$ and $\pi$ give $\cos\varphi_{\mathrm{oct}}=\pm 1$, which implies that the linearized LLG depends on which of the two saddle points are under consideration.
$\begin{pmatrix}\delta\dot{\theta}_{\mathrm{oct}}\\ \delta\dot{\varphi}_{\mathrm{oct}}\end{pmatrix}=\frac{\gamma}{1+\alpha^{2}}\begin{pmatrix}-\alpha H_{J}&\beta_{x}H_{\mathrm{K}}\\ -H_{J}&-\alpha\beta_{x}H_{\mathrm{K}}\end{pmatrix}\begin{pmatrix}\delta\dot{\theta}_{\mathrm{oct}}\\ \delta\dot{\varphi}_{\mathrm{oct}}\end{pmatrix},$ (26)
where $\beta_{x}=\mp\frac{H_{\mathrm{x,eff}}}{H_{\mathrm{K}}}-1$ for $\varphi_{\mathrm{oct}}=0$ and $\pi$, respectively. Therefore, the eigenvalue of the linearized LLG matrix is
$\lambda_{+}=\frac{\gamma}{(1+\alpha^{2})}\frac{-\alpha\left[H_{J}+\beta_{x}H_{K}\right]+\sqrt{\alpha^{2}\left(H_{J}-\beta_{x}H_{K}\right)^{2}-4\beta_{x}H_{J}H_{K}}}{2}.$ (27)
  • Eigenvalue of Hessian matrix
Because the saddle points are not identical, the escape processes must consider the two saddle points separately. For the saddle point at $\varphi_{\mathrm{oct}}=0$ and the minimum at $\cos\varphi_{\mathrm{oct}}=-\frac{H_{\mathrm{x,eff}}}{H_{\mathrm{K}}}$, the eigenvalues of the Hessian matrix are
$\varepsilon_{\mathrm{1,min}}$ $=\mathcal{\mathcal{V}}\frac{\partial^{2}\mathcal{H}_{\mathrm{oct}}}{\partial\theta_{\mathrm{oct}}^{2}}\Big{|}_{\theta_{\mathrm{oct}}=\pi/2,\varphi_{\mathrm{oct}}=\cos^{-1}\quantity(\frac{H_{\mathrm{x,eff}}}{H_{\mathrm{K}}})}=3M_{s}H_{J}\mathcal{V},$ (28a)
$\varepsilon_{\mathrm{1,sp}}$ $=\mathcal{V}\frac{\partial^{2}\mathcal{H}_{\mathrm{oct}}}{\partial\theta_{\mathrm{oct}}^{2}}\Big{|}_{\theta_{\mathrm{oct}}=\pi/2,\varphi_{\mathrm{oct}}=0}=3M_{s}H_{J}\mathcal{V},$ (28b)
$\varepsilon_{\mathrm{2,min}}$ $=\mathcal{V}\frac{\partial^{2}\mathcal{H}_{\mathrm{oct}}}{\partial\varphi_{\mathrm{oct}}^{2}}\Big{|}_{\theta_{\mathrm{oct}}=\pi/2,\varphi_{\mathrm{oct}}=\cos^{-1}\quantity(\frac{H_{\mathrm{x,eff}}}{H_{\mathrm{K}}})}=3M_{s}\mathcal{V}H_{\mathrm{K}}\quantity(1-\frac{H_{\mathrm{x,eff}}^{2}}{H_{\mathrm{K}}^{2}}),$ (28c)
$\varepsilon_{\mathrm{2,sp}}$ $=\mathcal{V}\frac{\partial^{2}\mathcal{H}_{\mathrm{oct}}}{\partial\varphi_{\mathrm{oct}}^{2}}\Big{|}_{\theta_{\mathrm{oct}}=\pi/2,\varphi_{\mathrm{oct}}=0}=-3M_{s}\mathcal{V}(H_{\mathrm{K}}+H_{\mathrm{x,eff}}),$ (28d)
which gives
$\sqrt{\frac{\prod_{j}\left|\varepsilon_{j,\text{sp}}\right|}{\prod_{j}\varepsilon_{j,\min}}}=\sqrt{\quantity|\frac{H_{\mathrm{K}}\quantity(H_{\mathrm{K}}+H_{\mathrm{x,eff}})}{H_{\mathrm{K}}^{2}-H_{\mathrm{x,eff}}^{2}}|}=\sqrt{\quantity|\frac{H_{\mathrm{K}}}{H_{\mathrm{K}}-H_{\mathrm{x,eff}}}|}.$ (29)
Similarly, at the saddle point $\varphi_{\mathrm{oct}}=\pi$ and minimum point $\varphi_{\mathrm{oct}}=-\frac{H_{\mathrm{x,eff}}}{H_{\mathrm{K}}}$, the eigenvalues of the Hessian matrix are
$\varepsilon_{\mathrm{1,min}}$ $=\mathcal{V}\frac{\partial^{2}\mathcal{H}_{\mathrm{oct}}}{\partial\theta_{\mathrm{oct}}^{2}}\Big{|}_{\theta_{\mathrm{oct}}=\pi/2,\varphi_{\mathrm{oct}}=\cos^{-1}\quantity(\frac{H_{\mathrm{x,eff}}}{H_{\mathrm{K}}})}=3M_{s}H_{J}\mathcal{V},$ (30a)
$\varepsilon_{\mathrm{1,sp}}$ $=\mathcal{V}\frac{\partial^{2}\mathcal{H}_{\mathrm{oct}}}{\partial\theta_{\mathrm{oct}}^{2}}\Big{|}_{\theta_{\mathrm{oct}}=\pi/2,\varphi_{\mathrm{oct}}=\pi}=3M_{s}H_{J}\mathcal{V},$ (30b)
$\varepsilon_{\mathrm{2,min}}$ $=\mathcal{V}\frac{\partial^{2}\mathcal{H}_{\mathrm{oct}}}{\partial\varphi_{\mathrm{oct}}^{2}}\Big{|}_{\theta_{\mathrm{oct}}=\pi/2,\varphi_{\mathrm{oct}}=\cos^{-1}\quantity(\frac{H_{\mathrm{x,eff}}}{H_{\mathrm{K}}})}=3M_{s}\mathcal{V}H_{\mathrm{K}}\quantity(1-\frac{H_{\mathrm{x,eff}}^{2}}{H_{\mathrm{K}}^{2}}),$ (30c)
$\varepsilon_{\mathrm{2,sp}}$ $=\mathcal{V}\frac{\partial^{2}\mathcal{H}_{\mathrm{oct}}}{\partial\varphi_{\mathrm{oct}}^{2}}\Big{|}_{\theta_{\mathrm{oct}}=\pi/2,\varphi_{\mathrm{oct}}=\pi}=-3M_{s}\mathcal{V}(H_{\mathrm{K}}-H_{\mathrm{x,eff}}),$ (30d)
which gives
$\sqrt{\frac{\prod_{j}\left|\varepsilon_{j,\text{sp}}\right|}{\prod_{j}\varepsilon_{j,\min}}}=\sqrt{\quantity|\frac{H_{\mathrm{K}}(H_{\mathrm{K}}+H_{\mathrm{x,eff}})}{H_{\mathrm{K}}^{2}-H_{\mathrm{x,eff}}^{2}}|}=\sqrt{\quantity|\frac{H_{\mathrm{K}}}{H_{\mathrm{K}}+H_{\mathrm{x,eff}}}|}.$ (31)
Therefore, the escape time with $H_{\mathrm{ax}}$-assisted switching at saddle point $\varphi_{\mathrm{oct}}=0$ is
$\tau_{\mathrm{escx,1}}=\quantity[A\quantity(\frac{\delta E}{k_{\mathrm{B}}T})]^{-1}\frac{4\pi(1+\alpha^{2})\exp(\Delta E_{\mathrm{x,k}}/k_{\mathrm{B}}T)\sqrt{H_{\mathrm{K}}}}{\gamma\sqrt{|H_{\mathrm{x,eff}}-H_{\mathrm{K}}|}\quantity(-\alpha\left[H_{J}+\beta_{x}H_{K}\right]+\sqrt{\alpha^{2}\left(H_{J}-\beta_{x}H_{K}\right)^{2}-4\beta_{x}H_{J}H_{K}})},$ (32)
and the escape time at the saddle point $\varphi_{\mathrm{oct}}=\pi$ is
$\tau_{\mathrm{escx,2}}=\quantity[A\quantity(\frac{\delta E}{k_{\mathrm{B}}T})]^{-1}\frac{4\pi(1+\alpha^{2})\exp(\Delta E_{\mathrm{x,k}}/k_{\mathrm{B}}T)\sqrt{H_{\mathrm{K}}}}{\gamma\sqrt{|H_{\mathrm{x,eff}}+H_{\mathrm{K}}|}\quantity(-\alpha\left[H_{J}+\beta_{x}H_{K}\right]+\sqrt{\alpha^{2}\left(H_{J}-\beta_{x}H_{K}\right)^{2}-4\beta_{x}H_{J}H_{K}})}.$ (33)
.2.3 Analysis of depopulation factor
To evaluate the depopulation factor, we define $\delta E=\delta\mathcal{F}_{\mathrm{oct}}\mathcal{V}$ where $\delta\mathcal{F}_{\mathrm{oct}}$ represents the energy density dissipated over an equal free energy contour containing the energy maximum. $\delta\mathcal{F}_{\mathrm{oct}}$ as a function of out-of-plane component $m_{\mathrm{oct,z}}$ and in-plane polar angle $\varphi_{\mathrm{oct}}$ is given as Rózsa et al. (2019); Konakanchi et al. (2025)
[…]
  • Depopulation factor with $H_{\mathrm{ay}}$
Figure 5: Deviation of the depopulation factor from unity as a function of $H_{\mathrm{ay}}$
The energy of the system, given by the the Hamiltonian at the saddle point$\Big{(}\theta_{\mathrm{oct}}=\pi/2,\varphi_{\mathrm{oct}}=\sin^{-1}(\frac{H_{\mathrm{y,eff}}}{H_{\mathrm{K}}})/\pi-\sin^{-1}(\frac{H_{\mathrm{y,eff}}}{H_{\mathrm{K}}})\Big{)}$, is given as  Rózsa et al. (2019); Konakanchi et al. (2025):
$\mathcal{H}_{\mathrm{oct}}(H_{\mathrm{ay}})$ $=\frac{3}{4}M_{\mathrm{s}}H_{\mathrm{K}}\cos(2\varphi_{\mathrm{oct}})+\frac{3}{2}M_{\mathrm{s}}H_{\mathrm{J}}m^{2}_{z}+3M_{\mathrm{s}}H_{\mathrm{y,eff}}\sin\varphi_{\mathrm{oct}}$ (35)
$=\frac{3}{4}M_{\mathrm{s}}H_{\mathrm{K}}-\frac{3}{2}M_{\mathrm{s}}H_{\mathrm{K}}\quantity(\frac{H_{\mathrm{y,eff}}}{H_{\mathrm{K}}})^{2}+3M_{\mathrm{s}}H_{\mathrm{y,eff}}\frac{H_{\mathrm{y,eff}}}{H_{\mathrm{K}}},$
and, therefore, the relationship between $m_{\mathrm{oct},z}$ and $\varphi_{\mathrm{oct}}$ is given as
$m_{\mathrm{oct},z}=\sqrt{\frac{H_{\mathrm{K}}^{2}\sin^{2}\varphi_{\mathrm{oct}}+H_{\mathrm{y,eff}}^{2}-2H_{\mathrm{K}}H_{\mathrm{y,eff}}\sin\varphi_{\mathrm{oct}}}{H_{\mathrm{J}}H_{\mathrm{K}}}}.$ (36)
Consequently, the integration in Eq. (34) is only dependent on $\varphi_{\mathrm{oct}}$, which gives
$\delta\mathcal{F}_{\mathrm{oct}}$ $=\int_{\pi/2}^{\pi}-3\alpha M_{\mathrm{s}}H_{J}\sqrt{\frac{H_{\mathrm{K}}^{2}\sin^{2}\varphi_{\mathrm{oct}}+H_{\mathrm{y,eff}}^{2}-2H_{\mathrm{K}}H_{\mathrm{y,eff}}\sin\varphi_{\mathrm{oct}}}{H_{\mathrm{J}}H_{\mathrm{K}}}}d\varphi_{\mathrm{oct}}$ (37)
×$\times\int_{\pi/2}^{\pi}\alpha(-3M_{\mathrm{s}}H_{\mathrm{K}}\sin\varphi_{\mathrm{oct}}\cos\varphi_{\mathrm{oct}}+3M_{\mathrm{s}}H_{\mathrm{y,eff}}\cos\varphi_{\mathrm{oct}})$
×$\times\sqrt{\frac{H_{\mathrm{J}}H_{\mathrm{K}}}{H_{\mathrm{K}}^{2}\sin^{2}\varphi_{\mathrm{oct}}+H_{\mathrm{y,eff}}^{2}-2H_{\mathrm{K}}H_{\mathrm{y,eff}}\sin\varphi_{\mathrm{oct}}}}\quantity(2H_{\mathrm{K}}^{2}\sin\varphi_{\mathrm{oct}}\cos\varphi_{\mathrm{oct}}-2H_{\mathrm{K}}H_{\mathrm{y,eff}}\cos\varphi_{\mathrm{oct}})d\varphi_{\mathrm{oct}}.$
Numerically integrating the above equation enables us to compute the inverse of the depopulation factor $\quantity[A\quantity(\frac{\delta E}{k_{\mathrm{B}}T})]^{-1}$. For the range of $H_{\mathrm{ay}}$ and $\mathcal{V}$ considered in this work, $\quantity[A\quantity(\frac{\delta E}{k_{\mathrm{B}}T})]^{-1}\approx 1$ with less than $5\%$ error, as presented in Fig. 5 and the error increases with $H_{\mathrm{ay}}$.
  • Depopulation factor with $H_{\mathrm{ax}}$
In an equal-free-energy contour including the saddle point ($\theta_{\mathrm{oct}}=\pi/2$, $\varphi_{\mathrm{oct}}=0$),
$\mathcal{H}_{\mathrm{oct}}(H_{\mathrm{ax}})$ $=\frac{3}{4}M_{\mathrm{s}}H_{\mathrm{K}}\cos(2\varphi_{\mathrm{oct}})+\frac{3}{2}M_{\mathrm{s}}H_{\mathrm{J}}m^{2}_{z}+3M_{\mathrm{s}}H_{\mathrm{x,eff}}\cos\varphi_{\mathrm{oct}}=\frac{3}{4}M_{\mathrm{s}}H_{\mathrm{K}}+3M_{\mathrm{s}}H_{\mathrm{x,eff}},$ (38)
Figure 6: Deviation of the depopulation factor from unity with $H_{\mathrm{ax}}$ for (a) $\varphi_{\mathrm{oct}}=0$ and (b) $\varphi_{\mathrm{oct}}=\pi$.
the relationship between $m_{\mathrm{oct},z}$ and $\varphi_{\mathrm{oct}}$ is given as
$m_{\mathrm{oct},z}=\sqrt{\frac{H_{\mathrm{K}}\sin^{2}\varphi_{\mathrm{oct}}+2H_{\mathrm{x,eff}}(1-\cos\varphi_{\mathrm{oct}})}{H_{\mathrm{J}}}}.$ (39)
Then Eq. (34) can be written as
$\delta\mathcal{F}_{\mathrm{oct}}$ $=\Bigg{|}\int_{\pi/2}^{0}-3\alpha M_{\mathrm{s}}H_{J}\sqrt{\frac{H_{\mathrm{K}}\sin^{2}\varphi_{\mathrm{oct}}+2H_{\mathrm{x,eff}}(1-\cos\varphi_{\mathrm{oct}})}{H_{\mathrm{J}}}}\mathrm{d}\varphi_{\mathrm{oct}}$ (40)
×$\times\int_{\pi/2}^{0}\alpha(-3M_{\mathrm{s}}H_{\mathrm{K}}\sin\varphi_{\mathrm{oct}}\cos\varphi_{\mathrm{oct}}-3M_{\mathrm{s}}H_{\mathrm{x,eff}}\sin\varphi_{\mathrm{oct}})$
×$\times\sqrt{\frac{H_{\mathrm{J}}}{H_{\mathrm{K}}\sin^{2}\varphi_{\mathrm{oct}}+2H_{\mathrm{x,eff}}(1-\cos\varphi_{\mathrm{oct}})}}(2H_{\mathrm{K}}\sin\varphi_{\mathrm{oct}}\cos\varphi_{\mathrm{oct}}+2H_{\mathrm{x,eff}}\sin\varphi_{\mathrm{oct}})\mathrm{d}\varphi_{\mathrm{oct}}\Bigg{|}.$
Numerically integrating the above equation allows us to compute $\quantity[A\quantity(\frac{\delta E}{k_{\mathrm{B}}T})]^{-1}$, which can again be approximated to $1$ with an error of less than $5\%$ for $\Delta E_{0}\geq 2k_{\mathrm{B}}T$, as shown in Fig. 6(a) and the error exponentially decreases as $\Delta E_{0}$ increases.
In then equal-free-energy contour including the saddle point ($\theta_{\mathrm{oct}}=\pi/2$, $\varphi_{\mathrm{oct}}=\pi$),
$\mathcal{H}_{\mathrm{oct}}$ $=\frac{3}{4}M_{\mathrm{s}}H_{\mathrm{K}}\cos(2\varphi_{\mathrm{oct}})+\frac{3}{2}M_{\mathrm{s}}H_{\mathrm{J}}m^{2}_{z}+3M_{\mathrm{s}}H_{\mathrm{x,eff}}\cos\varphi_{\mathrm{oct}}=\frac{3}{4}M_{\mathrm{s}}H_{\mathrm{K}}-3M_{\mathrm{s}}H_{\mathrm{x,eff}},$ (41)
$m_{\mathrm{oct},z}=\sqrt{\frac{H_{\mathrm{K}}\sin^{2}\varphi_{\mathrm{oct}}-2H_{\mathrm{x,eff}}(1+\cos\varphi_{\mathrm{oct}})}{H_{\mathrm{J}}}}.$ (42)
Then Eq. (34) can be written as
$\delta\mathcal{F}_{\mathrm{oct}}$ $=\Bigg{|}\int_{\pi/2}^{0}-3\alpha M_{\mathrm{s}}H_{J}\sqrt{\frac{H_{\mathrm{K}}\sin^{2}\varphi_{\mathrm{oct}}-2H_{\mathrm{x,eff}}(1+\cos\varphi_{\mathrm{oct}})}{H_{\mathrm{J}}}}\mathrm{d}\varphi_{\mathrm{oct}}$ (43)
×$\times\int_{\pi/2}^{0}\alpha(-3M_{\mathrm{s}}H_{\mathrm{K}}\sin\varphi_{\mathrm{oct}}\cos\varphi_{\mathrm{oct}}-3M_{\mathrm{s}}H_{\mathrm{x,eff}}\sin\varphi_{\mathrm{oct}})$
×$\times\sqrt{\frac{H_{\mathrm{J}}}{H_{\mathrm{K}}\sin^{2}\varphi_{\mathrm{oct}}-2H_{\mathrm{x,eff}}(1+\cos\varphi_{\mathrm{oct}})}}(2H_{\mathrm{K}}\sin\varphi_{\mathrm{oct}}\cos\varphi_{\mathrm{oct}}+2H_{\mathrm{x,eff}}\sin\varphi_{\mathrm{oct}})\mathrm{d}\varphi_{\mathrm{oct}}\Bigg{|}.$
Even in this case $\quantity[A\quantity(\frac{\delta E}{k_{\mathrm{B}}T})]^{-1}\approx 1$ with an error of less than $2\%$ when $\Delta E_{0}\geq 2k_{\mathrm{B}}T$, as shown in Fig. 6(b).
.3 Numerical Simulation
.3.1 Escape time definition
Figure 7: (a) Temporal evolution of $m_{\mathrm{oct,y}}$ and red circles highlight examples of back-and-forth fluctuations around $m_{\mathrm{oct,y}}=0$. (b) Field-free escape time as a function of $\Delta E_{0}$ with various $\xi$. (c) The rate of change of field-free $\tau_{\mathrm{esc}}$ with $\xi$.
The escape time from the up state to the down state is defined as the average time it takes to navigate from the up-state energy well to the down-state energy well. Each simulation is 1 ms long, and the ensemble size is chosen 8, which jointly ensures that at least $10^{3}$ inter-well switching events occur during each measurement. If the switching is defined corresponding to when $\mathbf{m}_{\mathrm{oct,y}}$ crosses the saddle point, as depicted in Fig. 7(a), it would over-estimate the number of switching events, thus underestimating the escape time. This is because $\mathbf{m}_{\mathrm{oct,y}}$ crosses the zero point multiple times within a single state switching event. For physically meaningful extraction, we introduce a parameter $\xi(>0)$, defined below, to obtain the escape time for the thermally activated process:
  • $\mathbf{m}_{\mathrm{oct}}$ switching from up state to down state is assumed to occur when $\mathbf{m}_{\mathrm{oct,y}}$ crosses $-\xi$,
  • $\mathbf{m}_{\mathrm{oct}}$ switching from down state to up state is assumed to occur when $\mathbf{m}_{\mathrm{oct,y}}$ crosses $\xi$.
Figure 7(b) shows the escape time of up state extracted from the LLG simulations with various $\xi$, while Fig. 7(c) shows the sensitivity of $\tau_{\mathrm{esc}}$ to $\xi$. In the field-free thermal fluctuations case, the escape time has a positive dependence on $\xi$, whereas the dependence shrinks as $\xi$ increases. We select $\xi=0.5$ as the criterion for switching because this ensures $\tau_{\mathrm{esc}}$ has a reduced dependence on $\xi$ and the result matches well with reported relaxation time Konakanchi et al. (2025) extracted from the magnetization relaxation process.
.3.2 Spin configuration
The unit cell of $\mathrm{Mn_{3}Sn}$ is composed of 6 $\mathrm{Mn}$ spins in two stacked kagome planes so the complete form of the Hamiltonian consisting 6 spins is Liu and Balents (2017):
$\mathcal{H}(\mathbf{m})$ $=J_{\mathrm{E}}((1+\delta_{\mathrm{E}})\mathbf{m}_{1}\cdot\mathbf{m}_{2}+\mathbf{m}_{2}\cdot\mathbf{m}_{3}+\mathbf{m}_{3}\cdot\mathbf{m}_{1})+J_{\mathrm{E}}((1+\delta_{\mathrm{E}})\mathbf{m}_{4}\cdot\mathbf{m}_{5}+\mathbf{m}_{5}\cdot\mathbf{m}_{6}+\mathbf{m}_{6}\cdot\mathbf{m}_{4}))$ (44)
$-J_{\mathrm{E}}(\mathbf{m}_{1}\cdot\mathbf{m}_{4}+\mathbf{m}_{2}\cdot\mathbf{m}_{6}+\mathbf{m}_{3}\cdot\mathbf{m}_{6})-\sum_{i=1}^{6}{\left[K_{\mathrm{u}}(\mathbf{m}_{i}\cdot\mathbf{u}_{i})^{2}+M_{\mathrm{s}}\mathbf{H}_{\mathrm{a}}\cdot\mathbf{m}_{i})\right]}$
$+D_{\mathrm{M}}\mathbf{z}\cdot(\mathbf{m}_{1}\times\mathbf{m}_{2}+\mathbf{m}_{2}\times\mathbf{m}_{3}+\mathbf{m}_{3}\times\mathbf{m}_{1})+D_{\mathrm{M}}\mathbf{z}\cdot(\mathbf{m}_{4}\times\mathbf{m}_{5}+\mathbf{m}_{5}\times\mathbf{m}_{6}+\mathbf{m}_{6}\times\mathbf{m}_{4}),$
where suffix $i=4,5,6$ represents another three sublattices in the second kagome plane. The Hamiltonian includes both intra-plane and inter-plane isotropic exchange interactions and the easy axes of these sublattices are $\mathbf{u}_{4}=\mathbf{u}_{1}=-1/2\mathbf{x}+\sqrt{3}/2\mathbf{y}$, $\mathbf{u}_{5}=\mathbf{u}_{2}=-1/2\mathbf{x}-\sqrt{3}/2\mathbf{y}$ and $\mathbf{u}_{6}=\mathbf{u}_{3}=\mathbf{x}$.
We further perform numerical simulations using both the 3-spin and 6-spin configurations. As shown in Fig. 8, the escape times from the two configurations overlap, indicating that the use of the 3-spin configuration in the main text is justified.
Figure 8: Escape time collected from 3-spin and 6-spin configurations with (a) $H_{\mathrm{ax}}$ and (b) $H_{\mathrm{ay}}$.
.3.3 Damping coefficient dependence
Figure 9: Field-assisted escape time with $\alpha$ ranging from $10^{-4}$ to $10^{-2}$ and $\Delta E_{0}=3k_{\mathrm{B}}T$
To investigate the intermediate-to-high damping regime and validate the applicable range of our theoretical model, we conducted numerical simulations over a broad range of damping coefficients $\alpha$ from $10^{-4}$ to $10^{-2}$. As shown in Fig. 9, the theoretical predictions for $H_{\mathrm{ax}}$-assisted escape remain valid across this entire range, whereas the theory fails for $H_{\mathrm{ay}}$ when $\alpha<5\times 10^{-3}$.
.3.4 Down-to-up escape time analysis
Figure 10: The schematic of escape time from down state to up state while numerical extracted data is in markers and analytical predictions are plotted with lines. $\Delta E_{0}$ ranges from $2k_{\mathrm{B}}T$ to $5k_{\mathrm{B}}T$ and field is applied along (a) $x$-axis and (b) $y$-axis.
According to the symmetry of the energy landscape, the escape times from down state to up state are $\tau_{\mathrm{escx,\downarrow}}=\tau_{\mathrm{escx,\uparrow}}$ and $\tau_{\mathrm{escy,\downarrow}}=\tau_{\mathrm{escy,\uparrow}}\exp((\Delta E_{\mathrm{y,2}}-\Delta E_{\mathrm{y,1}})k_{\mathrm{B}}T)$. Fig. 10 shows that our analytic model fits well with the numerical LLG data. It is worth mentioning that $\tau_{\mathrm{escy,\downarrow}}$ increases rapidly with $H_{\mathrm{ay}}$, so we restrict the maximum external field to $80~{}\mathrm{mT}$ for analysis (the field is restricted to $50~{}\mathrm{mT}$ with $\Delta E_{0}=5k_{\mathrm{B}}T$).